Jesse rents a moving van for $75 and must pay $2 per mile. The following week, Alex rents the same van, is charged $80 for the rental and $1.50 per mile. If they each paid the same amount and drove the same number of miles, how far did they each travel
Answer by mananth(16946)

number of miles x both same
75+2x=80+1.6x
2x-1.5x=80-75
0.5x= 5
x=5/0.5
x=10 miles